# FastDriveVLA: Efficient End-to-End Driving via Plug-and-Play Reconstruction-based Token Pruning

### v2

@@ Introduction
End-to-end autonomous driving (Hu et al. 2023; Prakash, Chitta, and Geiger 2021; Wu et al. 2022; Zhang et al. 2021; Jiang et al. 2023) has recently shown remarkable potential, promising to revolutionize future transportation systems. Unlike traditional modular autonomous driving systems—which divide the task into distinct components such as perception (Lang et al. 2019; Li et al. 2024), prediction (Gu, Sun, and Zhao 2021; Liu et al. 2021), and planning (Caesar et al. 2021; Ettinger et al. 2021)—end-to-end approaches learn the entire driving pipeline within a unified framework. This design not only mitigates error propagation between modules but also enhances system simplicity.
@@
Given the remarkable reasoning capabilities demonstrated by Vision-Language Models (VLMs) (Liu et al. 2023; Wang et al. 2024a; Cao et al. 2025) in visual question answering tasks, recent studies have explored their extension to embodied intelligence and autonomous driving by incorporating action generation capabilities. These models, referred to as Vision-Language-Action (VLA) models (Tian et al. 2024; Sima et al. 2024; Wang et al. 2025; Chi et al. 2025), are increasingly adopted in end-to-end autonomous driving systems and have demonstrated superior performance over traditional modular approaches. However, existing visual language models (VLMs) usually convert visual inputs into numerous visual tokens. This approach has also been adopted by visual language attention (VLA) models, leading to considerable computational overhead and increased inference latency. This presents a significant challenge for deploying vehicles in real-world scenarios, where both computational resources and inference speed are severely limited.
@@
Numerous efforts have been made to accelerate VLM inference by visual token reduction. Some approaches introduce newly designed multimodal projectors to compress visual tokens (Cha et al. 2024; Li et al. 2025; Hu et al. 2024; Cai et al. 2024; Zhang et al. 2025b), but these methods require retraining the entire model, making them computationally expensive. Other approaches attempt to remove redundant visual tokens in a plug-and-play manner (Shang et al. 2024; Yang et al. 2025b, a; Dhouib et al. 2025), which can be broadly categorized into attention-based (Chen et al. 2024; Zhang et al. 2024b; Xing et al. 2024; Zhao et al. 2025) and similarity-based (Wen et al. 2025; Zhang et al. 2024a; Alvar et al. 2025; Zhang et al. 2025a) pruning strategies. Attention-based methods depend significantly on precise text-vision alignment and are particularly vulnerable to irrelevant information in the visual tokens. This issue is further exacerbated in autonomous driving scenarios, where text inputs are typically fixed and concise, offering limited guidance for effective token selection. While similarity-based methods are also ill-suited for autonomous driving, where visual inputs often contain well-defined foreground regions, such as lanes, pedestrians, and vehicles. In such cases, emphasizing token similarity becomes less meaningful, and similarity-based pruning may mistakenly retain background tokens irrelevant to driving tasks.
@@
To facilitate the training of ReconPruner, we further introduce a large-scale dataset named nuScenes-FG. To construct this dataset, we first define the concept of foreground in autonomous driving scenes, and then leverage Grounded-SAM (Ren et al. 2024) to segment the nuScenes (Caesar et al. 2020) dataset accordingly. This large-scale dataset contains 241k image-mask pairs across six camera views, with segmentation annotations of foreground regions.
@@
End-to-End Autonomous Driving Research in autonomous driving has seen a notable evolution from conventional modular pipelines, which decompose the task into perception, prediction, and planning, towards unified end-to-end learning frameworks. Seminal work like PilotNet demonstrated the feasibility of directly mapping raw pixel inputs to vehicle control commands using deep neural networks (Bojarski et al. 2016). While early behavioral cloning methods demonstrated the promise of end-to-end driving, they suffered from critical issues such as causal confusion and covariate shift. A primary research thrust to mitigate these limitations involved injecting explicit guidance into the learning process; for example, Conditional Imitation Learning (CIL) (Codevilla et al. 2018) incorporated high-level navigational commands to regularize the driving policy. Concurrently, another line of work focused on enhancing model robustness through architectural innovation, with approaches like TransFuser (Prakash, Chitta, and Geiger 2021) leveraging Transformer architectures to effectively fuse multi-modal sensor data. More recently, works such as SOLVE (Wen et al. 2024) and OpenDriveVLA (Zeng et al. 2024) proposed to synergize the direct action generation of end-to-end networks with the power of large vision-language-action architectures to improve both interpretability and performance in complex scenarios.
@@
Driving Vision-Language-Action Models Recently, the integration of large language models (LLMs) has given rise to Vision-Language-Action (VLA) models, which are setting a new frontier in autonomous driving. These models aim to enhance the vehicle’s reasoning capabilities, interpretability, and ability to handle long-tail scenarios by grounding driving actions in natural language. DriveGPT4 (Xu et al. 2024) showcased how LLMs can be adapted for motion planning and vehicle control. Building upon this trend, OpenDriveVLA (Zeng et al. 2024) and Impromptu VLA (Sha et al. 2024) are significant contributions that focus on developing open-source, large-scale VLAs specifically for driving. They demonstrate how to train powerful models that can process complex visual scenes and generate fine-grained control actions. The development of such data-hungry models is critically dependent on comprehensive datasets. OmniDrive (Wang et al. 2024b) provides a holistic, vision-language dataset featuring rich annotations and counterfactual reasoning scenarios.
@@
Visual Token Pruning Existing VLMs convert visual inputs into a large number of tokens, leading to significant computational overhead and inference latency. Many studies have explored visual token pruning as a plug-and-play approach to improve the inference efficiency (Shang et al. 2024; Yang et al. 2025b; Xing et al. 2024; Wen et al. 2025; Zhang et al. 2025a), which can be broadly categorized based on their pruning criteria. The first category, attention-based methods, such as FastV (Chen et al. 2024) and SparseVLM (Zhang et al. 2024b), assesses the importance of visual tokens using attention scores from text tokens, which heavily rely on the correlation between user instructions and input images. However, in the driving tasks, where instructions are typically fixed and concise, this correlation is insufficient to guide effective token selection. The second category, similarity-based methods, like VisPruner (Zhang et al. 2024a) and DivPrune (Alvar et al. 2025), removes redundancy by selecting a diverse subset of visual tokens. Nevertheless, in the driving scenarios, input images often contain well-defined foreground regions, and excessive retention of background tokens irrelevant to the driving task can degrade performance under constrained computational budgets.
@@
The nuScenes (Caesar et al. 2020) dataset includes 3D bounding box annotations for humans and vehicles, yet this representation inherently captures extraneous background elements due to the coarse nature of axis-aligned bounding volumes. Although a subsequent map expansion package with 11 semantic layers is available, these annotations still fail to comprehensively cover all relevant regions. To address this, we employ Grounded-SAM (Ren et al. 2024) to generate consistent and fine-grained foreground segmentation annotations across nuScenes scenes. The resulting nuScenes-FG dataset comprises 241k image–mask pairs from six camera views, with examples shown in Fig. 2.
@@
We propose a novel and lightweight plug-and-play pruner named ReconPruner, which is trained via a pixel-level reconstruction. The architecture of ReconPruner consists of a PrunerLayer and a Scorer, as illustrated in Fig. 3. The PrunerLayer is implemented as a single decoder layer of Qwen2.5-VL-3B (Bai et al. 2025). The Scorer is implemented as a single-layer feedforward network with a weight shape of $\mathbb{R}^{D\times 1}$, where $D$ denotes the hidden state dimension. Overall, ReconPruner is highly lightweight, with a total size of only 0.07B parameters.
@@
Since our primary objective is to enable ReconPruner to effectively identify and select visual tokens that contain meaningful foreground information, we draw inspiration from prior masked image modeling (MIM) approaches (He et al. 2022; Xie et al. 2022) and design a MAE-style pixel reconstruction strategy. During training, we select the subset of visual tokens with the highest saliency scores as predicted by ReconPruner and use them for masked foreground reconstruction. The reconstruction loss computed on this subset serves as a supervisory signal, encouraging ReconPruner to assign higher saliency scores to visual tokens that genuinely correspond to foreground content.
@@
However, relying solely on foreground reconstruction can lead to a degenerate solution where ReconPruner takes a shortcut by indiscriminately assigning high saliency scores to all visual tokens, thus boosting the reconstruction performance. To address this issue, we draw inspiration from Generative Adversarial Networks (GANs) (Goodfellow et al. 2020) and propose an adversarial foreground-background reconstruction strategy. Specifically, ReconPruner is additionally required to reconstruct the background regions using the visual tokens that receive low saliency scores. By imposing this complementary background reconstruction constraint, the model is effectively discouraged from assigning uniformly high saliency scores, thereby promoting a more precise and discriminative scoring of visual tokens. This adversarial setup enhances ReconPruner’s ability to differentiate foreground tokens from background ones, resulting in improved token selection performance.
@@
where $M_{i}$ and $S_{i}$ denote the $i$-th element of $M$ and $S$, respectively. However, since $M$ is non-differentiable, directly masking visual tokens $V$ with mask $M$ would block the gradient flow during backpropagation. To address this issue, we adopt the Straight-Through Estimator (STE) (Bengio, Léonard, and Courville 2013) technique, which applies a discrete mask in the forward pass while using a continuous approximation in the backward pass to enable gradient propagation. This operation is defined as follows:
$\tilde{M}=M+\text{stop\_grad}(1-M),$ (4)
@@
The reconstruction decoder $D$ consists of six Qwen2.5-VL-3B (Bai et al. 2025) decoder layers and a feedforward reconstruction head. We feed both $V_{fore}$ and $V_{back}$ into the reconstruction decoder $D$ to obtain the reconstructed foreground $I^{pred}_{fore}\in\mathbb{R}^{3\times H\times W}$ and background $I^{pred}_{back}\in\mathbb{R}^{3\times H\times W}$, which can be formulated as follows:
@@
In order to leverage both pixel-level accuracy and perceptual consistency, we formulate the reconstruction loss as a weighted combination of the Mean Squared Error (MSE) and the Structural Similarity Index Measure (SSIM) loss (Wang et al. 2004), as defined below:
@@
[…]
$V_{\mathrm{select}}=\{v_{i}\mid i\in\mathcal{I}\},\quad\mathcal{I}=\text{TopK}(S,M).$ (9)
To ensure that the retained visual tokens preserve their original spatial semantics, we also retain their corresponding position embeddings. The selected visual tokens $V_{\mathrm{select}}\in\mathbb{R}^{M\times D}$ and the text tokens $T\in\mathbb{R}^{L\times D}$ are then jointly fed into the large language model $f_{\phi}$ to predict the final action, which can be formulated as:
$Action=f_{\phi}([V_{\mathrm{select}},T]).$ (10)
@@
Input size $\times 1596$ 1596, 3249 tokens ($\mathbf{100\%}$)
@@
We adopt Impromptu-VLA (Chi et al. 2025), the current state-of-the-art end-to-end VLA model for autonomous driving, as the base model for visual token pruning. It is built upon the Qwen2.5-VL (Bai et al. 2025) architecture. The encoder of Impromptu-VLA remains frozen during its original training process, making its parameters and architecture identical to those of Qwen2.5-VL. Since the reconstruction task is non-causal by nature, we replace the causal attention mechanism with full attention in both the ReconPruner and reconstruction decoder.
@@
We evaluate our method on the nuScenes (Caesar et al. 2020) dataset, a large-scale benchmark specifically designed for autonomous driving in urban environments. It consists of 1,000 driving scenes, each lasting approximately 20 seconds. For testing, we follow the official evaluation protocol of Impromptu-VLA and use a total of 6,019 test samples. Following prior work (Wang et al. 2025), we evaluate the performance of open-loop planning using three metrics: trajectory prediction L2 error, Collision Rate, and Intersection Rate with the road boundary.
@@
For comparison, we adopt FastV (Chen et al. 2024) and SparseVLM (Zhang et al. 2024b) as attention-based baselines, and DivPrune (Alvar et al. 2025) and VisPruner (Zhang et al. 2024a) as similarity-based baselines.
@@
As shown in Tab. 1, when pruning 25% of the visual tokens, our method outperforms all baseline methods across all metrics. Notably, our approach even surpasses the original unpruned model in terms of L2 and Collision metrics, with improvements of 0.1% and 1.0%, respectively. This encouraging result supports our hypothesis that focusing on foreground-relevant visual tokens is key to autonomous driving. When pruning 50% of the visual tokens, we observe an interesting phenomenon: most methods exhibit a better Collision performance compared to the 25% pruning setting. Similarly, at a 75% pruning ratio, some methods even achieve higher Intersection performance than at 50%. However, this performance improvement with increasing pruning ratios is not observed under the L2 metric. We attribute this to the relatively small absolute values of Collision and Intersection metrics, making them more susceptible to noise.
@@
Gt-mask$+$Text-attn 14.07 28.71 53.70 32.16 99.0% 0.00 0.16 0.63 0.26 92.4% 0.53 2.50 5.82 2.95 94.8%
@@
To achieve reconstruction-based visual token pruning, a straightforward approach is to resize the foreground mask to match the spatial resolution of the visual tokens and prune the tokens at the corresponding positions. However, this approach encounters two major challenges: (1) the foreground mask provides only binary cues and lacks the capacity to quantify the saliency of individual visual tokens, making it unsuitable for ranking and pruning at arbitrary ratios; and (2) the spatial alignment between the foreground mask and visual tokens is often inaccurate — prior work (Darcet et al. 2023) has shown that the positions of visual tokens generated by vision encoders frequently exhibit spatial misalignment with the original image patches.
To compare with the pruning method based on foreground masks, we use text attention to estimate the saliency of visual tokens and prioritize those located within the foreground mask region. We also compare this with a baseline that prunes solely based on text attention. As shown in Tab. 3, we find that pruning guided by foreground masks achieves a clear performance improvement over text-attention-only pruning, indicating that foreground visual tokens are indeed more informative. However, our method remains more efficient, as it addresses the spatial misalignment issue of foreground visual tokens. Moreover, generating foreground masks using Grounded-SAM (Ren et al. 2024) typically takes around 3 seconds per image, which incurs a prohibitive time cost for practical deployment.

### v4

@@ Introduction
End-to-end autonomous driving (hu2023planning; prakash2021multi; wu2022trajectory; zhang2021end; jiang2023vad) has recently shown remarkable potential, promising to revolutionize future transportation systems. Unlike traditional modular autonomous driving systems—which divide the task into distinct components such as perception (lang2019pointpillars; li2024bevformer), prediction (gu2021densetnt; liu2021multimodal), and planning (caesar2021nuplan; ettinger2021large)—end-to-end approaches learn the entire driving pipeline within a unified framework. This design not only mitigates error propagation between modules but also enhances system simplicity.
@@
Given the remarkable reasoning capabilities demonstrated by Vision-Language Models (VLMs) (liu2023visual; wang2024qwen2; cao2025move) in visual question answering tasks, recent studies have explored their extension to embodied intelligence and autonomous driving by incorporating action generation capabilities. These models, referred to as Vision-Language-Action (VLA) models (tian2024drivevlm; sima2024drivelm; wang2025omnidrive; li2025manipdreamer3d), are increasingly adopted in end-to-end autonomous driving systems and have demonstrated superior performance over traditional modular approaches. However, existing visual language models (VLMs) usually convert visual inputs into numerous visual tokens. This approach has also been adopted by visual language attention (VLA) models, leading to considerable computational overhead and increased inference latency. This presents a significant challenge for deploying vehicles in real-world scenarios, where both computational resources and inference speed are severely limited.
@@
Numerous efforts have been made to accelerate VLM inference by visual token reduction. Some approaches introduce newly designed multimodal projectors to compress visual tokens (cha2024honeybee; li2025tokenpacker; hu2024matryoshka; cai2024matryoshka; zhang2025llavamini), but these methods require retraining the entire model, making them computationally expensive. Other approaches attempt to remove redundant visual tokens in a plug-and-play manner (shang2024llavaprumerge; yang2025visionzip; yang2025topv; dhouib2025pact), which can be broadly categorized into attention-based (chen2024fastv; zhang2024sparsevlm; zhao2025gsearch) and similarity-based (zhang2024vispruner; alvar2025divprune; zhang2025cdpruner) pruning strategies. Attention-based methods depend significantly on precise text-vision alignment and are particularly vulnerable to irrelevant information in the visual tokens. This issue is further exacerbated in autonomous driving scenarios, where text inputs are typically fixed and concise, offering limited guidance for effective token selection. While similarity-based methods are also ill-suited for autonomous driving, where visual inputs often contain well-defined foreground regions, such as lanes, pedestrians, and vehicles. In such cases, emphasizing token similarity becomes less meaningful, and similarity-based pruning may mistakenly retain background tokens irrelevant to driving.
@@
To facilitate the training of ReconPruner, we further introduce a large-scale dataset named nuScenes-FG. To construct this dataset, we first define the concept of foreground in autonomous driving scenes, and then leverage Grounded-SAM (ren2024grounded) to segment the nuScenes (caesar2020nuscenes) dataset accordingly. This large-scale dataset contains 241k image-mask pairs across six camera views, with segmentation annotations of foreground regions.
@@
End-to-End Autonomous Driving Research in autonomous driving has seen a notable evolution from conventional modular pipelines, which decompose the task into perception, prediction, and planning, towards unified end-to-end learning frameworks. Seminal work like PilotNet demonstrated the feasibility of directly mapping raw pixel inputs to vehicle control commands using deep neural networks (bojarski2016end). While early behavioral cloning methods demonstrated the promise of end-to-end driving, they suffered from critical issues such as causal confusion and covariate shift. A primary research thrust to mitigate these limitations involved injecting explicit guidance into the learning process; for example, Conditional Imitation Learning (CIL) (codevilla2018imitation) incorporated high-level navigational commands to regularize the driving policy. Concurrently, another line of work focused on enhancing model robustness through architectural innovation, with approaches like TransFuser (prakash2021multi) leveraging Transformer architectures to effectively fuse multi-modal sensor data. More recently, works such as SOLVE (wen2024solve) and OpenDriveVLA (zeng2024opendrivevla) proposed to synergize the direct action generation of end-to-end networks with the power of large vision-language-action architectures to improve both interpretability and performance in complex scenarios.
@@
Driving Vision-Language-Action Models Recently, the integration of large language models (LLMs) has given rise to Vision-Language-Action (VLA) models, which are setting a new frontier in autonomous driving. These models aim to enhance the vehicle’s reasoning capabilities, interpretability, and ability to handle long-tail scenarios by grounding driving actions in natural language. DriveGPT4 (xu2024drivegpt4) showcased how LLMs can be adapted for motion planning and vehicle control. Building upon this trend, OpenDriveVLA (zeng2024opendrivevla) and Impromptu VLA (sha2024impromptu) are significant contributions that focus on developing open-source, large-scale VLAs specifically for driving. They demonstrate how to train powerful models that can process complex visual scenes and generate fine-grained control actions. The development of such data-hungry models is critically dependent on comprehensive datasets. OmniDrive (wang2024omnidrive) provides a holistic, vision-language dataset featuring rich annotations and counterfactual reasoning scenarios.
@@
Visual Token Pruning Existing VLMs convert visual inputs into a large number of tokens, leading to significant computational overhead and inference latency. Many studies have explored visual token pruning as a plug-and-play approach to improve the inference efficiency (shang2024llavaprumerge; yang2025visionzip; zhang2025cdpruner; chen2025streamkvstreamingvideoquestionanswering; ma2025mmg), which can be broadly categorized based on their pruning criteria. The first category, attention-based methods, such as FastV (chen2024fastv) and SparseVLM (zhang2024sparsevlm), assesses the importance of visual tokens using attention scores from text tokens, which heavily rely on the correlation between user instructions and input images. However, in the driving tasks, where instructions are typically fixed and concise, this correlation is insufficient to guide effective token selection. The second category, similarity-based methods, like VisPruner (zhang2024vispruner) and DivPrune (alvar2025divprune), removes redundancy by selecting a diverse subset of visual tokens. Nevertheless, in the driving scenarios, input images often contain well-defined foreground regions, and excessive retention of background tokens irrelevant to the driving task can degrade performance under constrained computational budgets.
@@
The nuScenes (caesar2020nuscenes) dataset includes 3D bounding box annotations for humans and vehicles, yet this representation inherently captures extraneous background elements due to the coarse nature of axis-aligned bounding volumes. Although a subsequent map expansion package with 11 semantic layers is available, these annotations still fail to comprehensively cover all relevant regions. To address this, we employ Grounded-SAM (ren2024grounded) to generate consistent and fine-grained foreground segmentation annotations across nuScenes scenes. The resulting nuScenes-FG dataset comprises 241k image–mask pairs from six camera views, with examples shown in Fig. 2.
@@
We propose a novel and lightweight plug-and-play pruner named ReconPruner, which is trained via a pixel-level reconstruction. The architecture of ReconPruner consists of a PrunerLayer and a Scorer, as illustrated in Fig. 3. The PrunerLayer is implemented as a single decoder layer of Qwen2.5-VL-3B (bai2025qwen2.5). The Scorer is implemented as a single-layer feedforward network with a weight shape of $\mathbb{R}^{D\times 1}$, where $D$ denotes the hidden state dimension. Overall, ReconPruner is highly lightweight, with a total size of only 0.07B parameters.
@@
Since our primary objective is to enable ReconPruner to effectively identify and select visual tokens that contain meaningful foreground information, we draw inspiration from prior masked image modeling (MIM) approaches (he2022masked; xie2022simmim) and design a MAE-style pixel reconstruction strategy. During training, we select the subset of visual tokens with the highest saliency scores as predicted by ReconPruner and use them for masked foreground reconstruction. The reconstruction loss computed on this subset serves as a supervisory signal, encouraging ReconPruner to assign higher saliency scores to visual tokens that genuinely correspond to foreground content.
@@
However, relying solely on foreground reconstruction can lead to a degenerate solution where ReconPruner takes a shortcut by indiscriminately assigning high saliency scores to all visual tokens, thus boosting the reconstruction performance. To address this issue, we draw inspiration from Generative Adversarial Networks (GANs) (goodfellow2020generative) and propose an adversarial foreground-background reconstruction strategy. Specifically, ReconPruner is additionally required to reconstruct the background regions using the visual tokens that receive low saliency scores. By imposing this complementary background reconstruction constraint, the model is effectively discouraged from assigning uniformly high saliency scores, thereby promoting a more precise and discriminative scoring of visual tokens. This adversarial setup enhances ReconPruner’s ability to differentiate foreground tokens from background ones, resulting in improved token selection performance.
@@
where $M_{i}$ and $S_{i}$ denote the $i$-th element of $M$ and $S$, respectively. However, since $M$ is non-differentiable, directly masking visual tokens $V$ with mask $M$ would block the gradient flow during backpropagation. To address this issue, we adopt the Straight-Through Estimator (STE) (bengio2013estimating) technique, which applies a discrete mask in the forward pass while using a continuous approximation in the backward pass to enable gradient propagation. This operation is defined as follows:
$\tilde{M}=S+\text{stop\_grad}(M-S),$ (4)
@@
The reconstruction decoder $D$ consists of six Qwen2.5-VL-3B (bai2025qwen2.5) decoder layers and a feedforward reconstruction head. We feed both $V_{fore}$ and $V_{back}$ into the reconstruction decoder $D$ to obtain the reconstructed foreground $I^{pred}_{fore}\in\mathbb{R}^{3\times H\times W}$ and background $I^{pred}_{back}\in\mathbb{R}^{3\times H\times W}$, which can be formulated as follows:
@@
In order to leverage both pixel-level accuracy and perceptual consistency, we formulate the reconstruction loss as a weighted combination of the Mean Squared Error (MSE) and the Structural Similarity Index Measure (SSIM) loss (wang2004image), as defined below:
@@
During inference, ReconPruner assigns saliency scores $S$ to a sequence of $N$ visual tokens. Given a target pruning ratio $p\in[0,1]$, we apply a Top-$K$ selection strategy to retain the top $K=\lfloor N\cdot(1-p)\rfloor$ visual tokens with the highest saliency scores, which can be formulated as:
$V_{\mathrm{select}}=\{v_{i}\mid i\in\mathcal{I}\},\quad\mathcal{I}=\text{TopK}(S,K).\$(9)
To ensure that the retained visual tokens preserve their original spatial semantics, we also retain their corresponding position embeddings. The selected visual tokens $V_{\mathrm{select}}\in\mathbb{R}^{K\times D}$ and the text tokens $T\in\mathbb{R}^{L\times D}$ are then jointly fed into the large language model $f_{\phi}$ to predict the final action, which can be formulated as:
$Action=f_{\phi}([V_{\mathrm{select}},T]).\$(10)
@@
Input size $1596\times 1596$, 3249 tokens ($\mathbf{100\%}$)
@@
We adopt Impromptu-VLA (chi2025impromptu), the current state-of-the-art end-to-end VLA model for autonomous driving, as the base model for visual token pruning. It is built upon the Qwen2.5-VL (bai2025qwen2.5) architecture. The encoder of Impromptu-VLA remains frozen during its original training process, making its parameters and architecture identical to those of Qwen2.5-VL. Since the reconstruction task is non-causal by nature, we replace the causal attention mechanism with full attention in both the ReconPruner and reconstruction decoder.
@@
We evaluate our method on the nuScenes (caesar2020nuscenes) dataset, a large-scale benchmark specifically designed for autonomous driving in urban environments. It consists of 1,000 driving scenes, each lasting approximately 20 seconds. For testing, we follow the official evaluation protocol of Impromptu-VLA and use a total of 6,019 test samples. Following prior work (wang2025omnidrive), we evaluate the performance of open-loop planning using three metrics: trajectory prediction L2 error, Collision Rate, and Intersection Rate with the road boundary.
@@
For comparison, we adopt FastV (chen2024fastv) and SparseVLM (zhang2024sparsevlm) as attention-based baselines, and DivPrune (alvar2025divprune) and VisPruner (zhang2024vispruner) as similarity-based baselines.
@@
As shown in Tab. 1, when pruning 25% of the visual tokens, our method outperforms all baseline methods across all metrics. Notably, our approach even surpasses the original unpruned model in terms of L2 and Intersection metrics, with improvements of 0.1% and 1.0%, respectively. This encouraging result supports our hypothesis that focusing on foreground-relevant visual tokens is key to autonomous driving. When pruning 50% of the visual tokens, we observe an interesting phenomenon: most methods exhibit a better Collision performance compared to the 25% pruning setting. Similarly, at a 75% pruning ratio, some methods even achieve higher Intersection performance than at 50%. However, this performance improvement with increasing pruning ratios is not observed under the L2 metric. We attribute this to the relatively small absolute values of Collision and Intersection metrics, making them more susceptible to noise.
@@
Gt-mask++Text-attn 14.07 28.71 53.70 32.16 99.0% 0.00 0.16 0.63 0.26 92.4% 0.53 2.50 5.82 2.95 94.8%
@@
To achieve reconstruction-based visual token pruning, a straightforward approach is to resize the foreground mask to match the spatial resolution of the visual tokens and prune the tokens at the corresponding positions. However, this approach encounters two major challenges: (1) the foreground mask provides only binary cues and lacks the capacity to quantify the saliency of individual visual tokens, making it unsuitable for ranking and pruning at arbitrary ratios; and (2) the spatial alignment between the foreground mask and visual tokens is often inaccurate — prior work (darcet2023vision) has shown that the positions of visual tokens generated by vision encoders frequently exhibit spatial misalignment with the original image patches.
To compare with the pruning method based on foreground masks, we use text attention to estimate the saliency of visual tokens and prioritize those located within the foreground mask region. We also compare this with a baseline that prunes solely based on text attention. As shown in Tab. 3, we find that pruning guided by foreground masks achieves a clear performance improvement over text-attention-only pruning, indicating that foreground visual tokens are indeed more informative. However, our method remains more efficient, as it addresses the spatial misalignment issue of foreground visual tokens. Moreover, generating foreground masks using Grounded-SAM (ren2024grounded) typically takes around 3 seconds per image, which incurs a prohibitive time cost for practical deployment.